# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v6

@@ A.4 Sufficiency for differentiability
This form of the scaling equation derived in the previous section is rather interesting. Since the ratio of two successive scaling terms is $(1-a_{i+1})a_{i+2}$, factors of both $a_{i}$ and $1-a_{i}$ are present in $s_{i}$. This has a few implications, firstly if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth. Secondly, having both $a_{i}$ and $1-a_{i}$ will cancel whichever slope multiplier $T_{i}$ contributes to $W_{i}$ at each point $x$, leaving behind the other term, which is less than 1.
@@
We begin by considering Equation 5 for layer $i$ (it equals 0 by Theorem A.3).
$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
We will prove our result by substituting Equation 4 into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each occurrence of $s$ in terms of $s_{i}$. Equation 4 gives a recurrence relation. Converting it to an non-recursive representation we have:
$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (7)
@@
$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (8)
This equation has a meaningful interpretation that is important to the argument. 1 is the full size of the initial derivative discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added. Every time a wave is added, it subtracts the first term appearing on the right hand side. The following argument shows that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining discontinuity, guaranteeing its disappearance in the limit. Inductively we can show:
$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (9)
In words this means that as the first term appearing on the right in Equation 8 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$
$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
@@
$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
@@ A.5 Error Decay
$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
@@ A.6 Second Derivatives
Let $R$ and $L$ denote the magnitude of $W^{\prime}_{i}$ on the left and right of $x$. $x$ is a peak location of $W_{i}$, so the right side slope is negative and the left is positive. Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})/L$ and $r_{1}=x+(1-a_{i+1})/R$.
On each subsequent iteration $i+n$ ($n\geq 2$), $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$. Since $x$ is a valley point the right slope is positive and the left is negative. The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $W_{i+n}$ ranges from 0 to 1 over these spans. Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$. The resulting non-recursive formulas are:
$x-l_{n}=\frac{1-a_{i+1}}{L}\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{1-a_{i+1}}{R}\prod_{m=2}^{n}a_{i+m}$ (10)
@@
The points $l_{n}$ and $r_{n}$ are all peak locations, Equation 6 gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$. Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $Rs_{i}$ rather than 1. Focusing on the right hand side we get:
$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=R*s_{i}\prod_{m=2}^{n}(1-a_{i+m})$
@@
$\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
The issue which arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$. The only way for these formulas to agree then is for $R^{2}=L^{2}$ which implies $a_{i}=1-a_{i}=0.5$. Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$). ∎
@@ A.7 Monotonicity and Continuity of Derivatives
Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure10). This can extend into the limit to show that $F^{\prime}$ is monotone decreasing and that $F$ is concave.
@@
To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i+1}[2^{i}]]$ for $0\leq n\leq 2^{i}$, which tracks the values of $F^{\prime}$ at the $i$th set of valley points. All but the first and last points will have been peaks at some point in their history, so Equation 6 gives the value of those derivatives as $f^{\prime}_{i}$.
We establish two inductive invariants. One is that the y-values in the list $Y_{i}$ remain sorted in descending order. The other is that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$, indicating that the constant value segments of $f_{i}$ lie in between the limits in the list $Y_{i}$. Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave. Which then be used to prove that their limit $F$ is also concave.
As the base case $f_{0}$ is a line with derivative 0, $V_{0}$ contains its two endpoints. $Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore both the points in $Y^{\prime}$ are in descending sorted order. The second part of the invariant is true since 0 is in between those values.
Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value, which we know from the second inductive hypothesis is in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$. The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$. This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order of $Y^{\prime}$.
@@
We will now show by contradiction that the limit $F$ of the sequence of concave $f_{i}$ is also concave. Assume that $F$ is non-concave. Then there exist points $a$, $b$, and $c$ such that $F(b)$ lies strictly below the line connecting the points $(a,F(a))$ and $(c,F(c))$. Lets say it’s below the line by an amount $\epsilon$. Since at each point $f_{i}$ converges to $F$, we can find $i_{a}$ such that $f_{ia}(a)-F(a)<\epsilon/2$, etc…, we take $i=\max(i_{a},i_{b},i_{c})$. Since $f_{i}(a)$ and $f_{i}(c)$ are no more than $\epsilon/2$ lower than their limiting values, the entire line connecting $(a,f_{i}(a))$ and $(c,f_{i}(c))$ is no more than $\epsilon/2$ lower than the line between $(a,F(a))$ and $(c,F(c))$. $f_{i}(b)$ is also no more than $\epsilon/2$ higher than $F(b)$, thus $f_{i}(b)$ must still lie below the line between $(a,f_{i}(a))$ and $(c,f_{i}(c))$, making $f_{i}$ non-concave and producing a contradiction. ∎
Lastly we briefly sketch out why $F^{\prime}$ is continuous. It relies on some of our earlier reasoning. Monotonicity of the derivative makes continuity easy to show, because when $x_{1}$ within $\delta$ of $x_{2}$ has $f(x_{1})$ within $\epsilon$ of $f(x_{2})$, so do all intermediate values of $x$. We can establish continuity of $F^{\prime}$ at bend points $x$ easily by using Equation 9. We can pick an $i$ so that the constant value segments of $f^{\prime}_{i}$ are within $\epsilon$ of $F^{\prime}(x)$ and then use the next iteration of bend points (since the constant intervals split, but the new segments near $x_{n}$ converge monotonically towards it) to find $\delta$. In the case of continuity for non bend points $x$, they sit inside a constant-valued interval of $f_{i}$ for each $i$. We can choose $i$ such that $\sum_{n=i}^{\infty}(1-c)^{n}<\epsilon/2$ because this series constrains how far derivative values can move in the limit, and then use the constant interval $x$ is situated in to find $\delta$.
@@ A.8 Activation Region Counting and Zaslavsky’s Theorem
In the coming sections, we present the results of using our 4-neuron-wide construction as an activation function on more complicated datasets. It will be useful to know the impact of this approach on the activation region count when assessing how to best spend a network’s parameter budget. To build up to it, we will provide an informal introduction of Zaslavsky’s theorem (Zaslavsky, 1975). The activation boundaries of ReLU neurons in a single hidden layer network form a hyperplane arrangement. Counting the number of cells, which are convex connected components of space separated by the hyperplanes, is equivalent to determining the number of activation regions produced by the shallow ReLU network. A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc… General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network). Zaslavsky’s theorem states that
@@
$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}$
To compute this sum, one would go down to the row of Pascal’s triangle corresponding to the number of planes $n$, and sum the first $D$ entries (see Figure 12 for a visual example). It’s a known fact that the sequence of the $d$th entry for each row grows to the $d$th power. So as the number of hyperplanes tends towards infinity, this sum is dominated by the $D$th term, which grows to the power $D$. In other words, the number of linear regions formed by a single hidden layer ReLU network grows exponentially with respect to its input dimension. This property partly explains why neural networks are successful at learning high dimensional functions, even when other classical methods become intractable.
@@
Figure 12: 3 lines dividing 2-dimensional space creates 7 cells. This is calculated by going down to row 3 (for the number of lines) and summing the numbers at indices 0, 1, and 2 (for the dimension). Also note that there is one 2-D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
@@
Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e. points, lines, planes, etc…) are given by the entries in Pascal’s triangle. So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower dimensional objects, or that each lower dimensional object could be ‘assigned’ a cell. There’s a striking similarity in the intuitions behind the recurrence that generates pascal’s triangle, and the calculation of the numbers of lower dimensional objects created from a hyperplane arrangement. In the case of Pascal’s triangle, if one wished to know how to choose 3 items from 6, one could reason that if item 6 is included, 2 items would be chosen from the remaining 5, and if item 6 is excluded, 3 would be chosen from the remaining 5, and therefore $\binom{6}{3}=\binom{5}{2}+\binom{5}{3}$. Now consider the case of 3 planes partitioning 3-space, and the addition of a 4th plane. The total number of lower dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points). One could also picture that the new 4th plane ‘cuts’ the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 12), and count the 2-cells that appear. Either reasoning leads to $R\binom{n}{D}=R\binom{n-1}{D}+R\binom{n-1}{D-1}$. This similarity between the recurrences means that if the initial conditions can be lined up suitably, Zaslavsky’s theorem can be proven by induction over Pascal’s triangle.
@@
In the case of our architecture, using our 1-dimensional construction as an activation function will generate $n$ generally configured sets of $m$ parallel hyperplanes. To account for this, a slight modification is made to the recurrence. Adding an additional set of $m$ hyperplanes would add $m$ times as many regions, as each of the lower dimensional objects in the existing configuration would project onto each of the $m$ additional hyperplanes. The effect that this has on the formula is that $R\binom{n,m}{D}=R\binom{n-1,m}{D}+m\times R\binom{n-1,m}{D-1}$. Inducting over Pascal’s triangle (with a base case of $R\binom{n,m}{1}=1+nm$) would give the non-recursive formula of $R\binom{n,m}{D}=\sum_{d=0}^{D}m^{d}\binom{n}{d}$. This means that out architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic trade off once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations). Of course, the effect on accuracy will all depend on how well these regions can be utilized.
A.9 Learning Rates

### v7

@@ A.4 Sufficiency for differentiability
This form of the scaling equation derived in the previous section is rather interesting. Since the ratio of two successive scaling terms is $(1-a_{i+1})a_{i+2}$, factors of both $a_{i}$ and $1-a_{i}$ are present in $s_{i}$. This has a few implications. Firstly, if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth. Secondly, having both $a_{i}$ and $1-a_{i}$ will cancel whichever slope multiplier $T_{i}$ contributes to $W_{i}$ at each point $x$, leaving behind the other term, which is less than 1.
@@
We begin by considering Equation 4 for layer $i$ (it equals 0 by Theorem A.3).:
$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$
We will prove our result by substituting Equation 3 into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each occurrence of $s$ in terms of $s_{i}$. Equation 3 gives a recurrence relation. Converting it to an non-recursive representation, we have:
$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right).$ (6)
@@
$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}).$ (7)
This equation has a meaningful interpretation that is important to the argument. 1 is the full size of the initial derivative discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added. Every time a wave is added, it subtracts the first term appearing on the right hand side. The following argument shows that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining discontinuity, guaranteeing its disappearance in the limit. Inductively, we can show:
$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m}).$ (8)
In words, this means that as the first term appearing on the right in Equation 7 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$,
$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m})),$
@@
$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m})).$
@@ A.5 Error Decay
$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}.$
To maximize $s_{i+2}$, we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
@@ A.6 Second Derivatives
Let $R$ and $L$ denote the magnitude of $W^{\prime}_{i}$ on the left and right of $x$. $x$ is a peak location of $W_{i}$, so the right side slope is negative and the left is positive. Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side gives $l_{1}=x-(1-a_{i+1})/L$ and $r_{1}=x+(1-a_{i+1})/R$.
On each subsequent iteration $i+n$ ($n\geq 2$), $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$. Since $x$ is a valley point, the right slope is positive and the left is negative. The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $W_{i+n}$ ranges from 0 to 1 over these spans. Solving for the new peaks again gives $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$. The resulting non-recursive formulas are:
$x-l_{n}=\frac{1-a_{i+1}}{L}\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{1-a_{i+1}}{R}\prod_{m=2}^{n}a_{i+m}.$ (9)
@@
The points $l_{n}$ and $r_{n}$ are all peak locations, Equation 5 gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$. Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 8 with the initial discontinuity size set to $Rs_{i}$ rather than 1. Focusing on the right hand side, we get:
$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=R*s_{i}\prod_{m=2}^{n}(1-a_{i+m}).$
@@
$\frac{R^{2}s_{i}}{(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}.$
The issue that arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$. The only way for these formulas to agree, then, is for $R^{2}=L^{2}$, which implies $a_{i}=1-a_{i}=0.5$. Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$). ∎
@@ A.7 Monotonicity and Continuity of Derivatives
Each of the $f^{\prime}_{i}$ are composed of constant value segments. We will show that those values are monotonically decreasing (this can be seen in Figure11). This can extend into the limit to show that $F^{\prime}$ is monotone decreasing and that $F$ is concave.
@@
To establish this, we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i+1}[2^{i}]]$ for $0\leq n\leq 2^{i}$, which tracks the values of $F^{\prime}$ at the $i$th set of valley points. All but the first and last points will have been peaks at some point in their history, so Equation 5 gives the value of those derivatives as $f^{\prime}_{i}$.
We establish two inductive invariants. One is that the y-values in the list $Y_{i}$ remain sorted in descending order. The other is that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$, indicating that the constant value segments of $f_{i}$ lie in between the limits in the list $Y_{i}$. Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave. This can then be used to prove that their limit $F$ is also concave.
As the base case, $f_{0}$ is a line with derivative 0, and $V_{0}$ contains its two endpoints. $Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes. Therefore, both the points in $Y^{\prime}$ are in descending sorted order. The second part of the invariant is true since 0 is in between those values.
Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$. This entire interval is between two valley points, so $f^{\prime}_{i}$ (which has not added $W_{i}$ yet) is some constant value, which we know from the second inductive hypothesis is in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$. The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$. This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order of $Y^{\prime}$.
@@
We will now show by contradiction that the limit $F$ of the sequence of concave $f_{i}$ is also concave. Assume that $F$ is non-concave. Then there exist points $a$, $b$, and $c$ such that $F(b)$ lies strictly below the line connecting the points $(a,F(a))$ and $(c,F(c))$. Let us imagine it is below the line by an amount $\epsilon$. Since at each point $f_{i}$ converges to $F$, we can find $i_{a}$ such that $f_{ia}(a)-F(a)<\epsilon/2$, etc…, we take $i=\max(i_{a},i_{b},i_{c})$. Since $f_{i}(a)$ and $f_{i}(c)$ are no more than $\epsilon/2$ lower than their limiting values, the entire line connecting $(a,f_{i}(a))$ and $(c,f_{i}(c))$ is no more than $\epsilon/2$ lower than the line between $(a,F(a))$ and $(c,F(c))$. $f_{i}(b)$ is also no more than $\epsilon/2$ higher than $F(b)$, thus $f_{i}(b)$ must still lie below the line between $(a,f_{i}(a))$ and $(c,f_{i}(c))$, making $f_{i}$ non-concave and producing a contradiction. ∎
Lastly, we briefly sketch out why $F^{\prime}$ is continuous. It relies on some of our earlier reasoning. Monotonicity of the derivative makes continuity easy to show, because when $x_{1}$ within $\delta$ of $x_{2}$ has $f(x_{1})$ within $\epsilon$ of $f(x_{2})$, so do all intermediate values of $x$. We can establish continuity of $F^{\prime}$ at bend points $x$ easily by using Equation 8. We can pick an $i$ so that the constant value segments of $f^{\prime}_{i}$ are within $\epsilon$ of $F^{\prime}(x)$ and then use the next iteration of bend points (since the constant intervals split, but the new segments near $x_{n}$ converge monotonically towards it) to find $\delta$. In the case of continuity for non bend points $x$, they sit inside a constant-valued interval of $f_{i}$ for each $i$. We can choose $i$ such that $\sum_{n=i}^{\infty}(1-c)^{n}<\epsilon/2$ because this series constrains how far derivative values can move in the limit, and then use the constant interval $x$ is situated in to find $\delta$.
@@ A.8 Activation Region Counting and Zaslavsky’s Theorem
In the coming sections, we present the results of using our 4-neuron-wide construction as an activation function on more complicated datasets. It will be useful to know the impact of this approach on the activation region count when assessing how to best spend a network’s parameter budget. To build up to it, we will provide an informal introduction of Zaslavsky’s theorem (Zaslavsky, 1975). The activation boundaries of ReLU neurons in a single hidden layer network form a hyperplane arrangement. Counting the number of cells, which are convex connected components of space separated by the hyperplanes, is equivalent to determining the number of activation regions produced by the shallow ReLU network. A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc. General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network). Zaslavsky’s theorem states that
@@
$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}.$
[…]
@@
Figure 13: 3 lines dividing 2-dimensional space creates 7 cells. This is calculated by going down to row 3 (for the number of lines) and summing the numbers at indices 0, 1, and 2 (for the dimension). Also note that there is one 2D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
@@
Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e., points, lines, planes, etc.) are given by the entries in Pascal’s triangle. So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower-dimensional objects, or that each lower-dimensional object could be “assigned” a cell. There is a striking similarity in the intuitions behind the recurrence that generates Pascal’s triangle, and the calculation of the numbers of lower-dimensional objects created from a hyperplane arrangement. In the case of Pascal’s triangle, if one wished to know how to choose 3 items from 6, one could reason that if item 6 is included, 2 items would be chosen from the remaining 5, and if item 6 is excluded, 3 would be chosen from the remaining 5, and therefore $\binom{6}{3}=\binom{5}{2}+\binom{5}{3}$. Now consider the case of 3 planes partitioning 3-space, and the addition of a 4th plane. The total number of lower-dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points). One could also picture that the new 4th plane “cuts” the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 13), and count the 2-cells that appear. Either reasoning leads to $R\binom{n}{D}=R\binom{n-1}{D}+R\binom{n-1}{D-1}$. This similarity between the recurrences means that if the initial conditions can be lined up suitably, Zaslavsky’s theorem can be proven by induction over Pascal’s triangle.
@@
In the case of our architecture, using our 1-dimensional construction as an activation function will generate $n$ generally configured sets of $m$ parallel hyperplanes. To account for this, a slight modification is made to the recurrence. Adding an additional set of $m$ hyperplanes would add $m$ times as many regions, as each of the lower dimensional objects in the existing configuration would project onto each of the $m$ additional hyperplanes. The effect that this has on the formula is that $R\binom{n,m}{D}=R\binom{n-1,m}{D}+m\times R\binom{n-1,m}{D-1}$. Inducting over Pascal’s triangle (with a base case of $R\binom{n,m}{1}=1+nm$) would give the non-recursive formula of $R\binom{n,m}{D}=\sum_{d=0}^{D}m^{d}\binom{n}{d}$. This means that our architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic tradeoff once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations). Of course, the effect on accuracy will all depend on how well these regions can be utilized.
Appendix B Additional Experiments
B.1 Learning Rates